OpenStudy (anonymous):

In 1492, Christopher Columbus journeyed to the New World, where he explored portions of the coast of two large Caribbean islands, Cuba and Hispaniola. Cuba is the larger of the two islands. Cuba is 10,000 square miles larger than Hispaniola. Hispaniola is only as large as Cuba. The equation below, where x is the area of Cuba in square miles, expresses this relationship. x − 10,000 = 3/4x What is the area of Cuba, in square miles?

OpenStudy (anonymous):

ummm

OpenStudy (anonymous):

1/4 x=10,000

OpenStudy (anonymous):

1/40,000

OpenStudy (anonymous):

this is confuseing

OpenStudy (anonymous):

was i right caz i need the answeer to be check im rather in a hurry

OpenStudy (anonymous):

umm i just guess caz ill never get it

OpenStudy (anonymous):

2,500

OpenStudy (anonymous):

oh so the answer is x=2,500

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2,500*4=10,000

OpenStudy (anonymous):

i swear retricehole go away caz i now how to and i did this is whhy im here not like u to put people don go away

OpenStudy (anonymous):

i never asked for your help ok i ask people that are pateinet can actually do this

OpenStudy (anonymous):

Alright. As Hero was saying, when you have \[\frac{ a }{ b } = \frac{ c }{ d }\] then you can make is so that it's ad = bc.

OpenStudy (anonymous):

This is called cross-multiplying, because it looks like this: